OpenStudy (anonymous):

The area is the product of the tablet length, 22+2x, and the tablet width, 17+2x. This product equals 374+78+4x^2 I want to know how the got that.

OpenStudy (anonymous):

@hoblos can you help?

OpenStudy (anonymous):

all i need to know how they got the 374+78+4x^2

OpenStudy (hoblos):



OpenStudy (hoblos):

374+78+4x^2 or 374+78x+4x^2

OpenStudy (anonymous):

oh yea my bad

OpenStudy (anonymous):

374+78x+4x^2

OpenStudy (hoblos):

ok so we have to use (a+b)(c+d) = a(c+d) + b(c+d)

OpenStudy (hoblos):

for a(c+d) , we have to distribute a into (c+d) so we get ac + ad

OpenStudy (hoblos):

so (22+2x)(17+2x) = 22(17+2x) + 2x(17+2x) = 22(17) + 22(2x) + 2x(17) + 2x(2x) what do you get if you simplify now ?

OpenStudy (anonymous):

374+44x+34x+4x^2

OpenStudy (anonymous):

oh ok thanks

OpenStudy (hoblos):

any time :)